OpenStudy (anonymous):

The ratio of the perimeters of two similar triangles is 2:3. What is the area of each of these triangles if the sum of their area is 260cm^?

OpenStudy (callisto):

Why do you post the same question twice? If you don't understand, ask what you don't understand. Making the same post twice gives an impression that you're not respecting those who answered your questions, especially when you didn't ask what you don't understand.

OpenStudy (anonymous):

@Callisto I dont get what to do next

OpenStudy (anonymous):

I've go the equation a+260-a=260

OpenStudy (callisto):

Can you express A2 in terms of A1?

OpenStudy (callisto):

Using the property \[(\frac{P_1}{P_2})^2=\frac{A_1}{A_2}\]

OpenStudy (anonymous):

A1=260-A2

OpenStudy (callisto):

Try again using that property. \(\frac{P_1}{P_2}\) is the ratio of the perimeters.

OpenStudy (anonymous):

A1=(A2)(P1/P2)^2

OpenStudy (callisto):

Okay. Next, can you replace (p1/P2) by the numbers given in the question? Please refer to my previous comment or ask if you're lost!